Company manufactures one product that goes through one processing department called Mixing. All raw materials are introduced at the start of work in the Mixing Department. The company uses the weighted-average method of process costing. Its Work in Process T-account for the Mixing Department for June follows (all forthcoming questions pertain to June):

Work in Process—Mixing Department
June 1 balance

25,000

Completed and transferred
to Finished Goods
?
Materials 157,080
Direct labor 99,500
Overhead 117,000
June 30 balance ?

The June 1 work in process inventory consisted of 4,000 units with $13,020 in materials cost and $11,980 in conversion cost. The June 1 work in process inventory was 100% complete with respect to materials and 60% complete with respect to conversion. During June, 36,500 units were started into production. The June 30 work in process inventory consisted of 9,600 units that were 100% complete with respect to materials and 50% complete with respect to conversion.

1. Prepare the journal entry to record the overhead cost applied to production. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)

2. How many units were completed and transferred to finished goods during the period?

3. Compute the equivalent units of production for materials.
Answer #1

1) Journal entry

Date account and explanation Debit Credit
Work in process 117000
Manufacturing overhead 117000
(To record overhead applied)

2) Unit completed and transferred out = 4000+36500-9600 = 30900

3) Equivalent unit of material = 30900+9600*100% = 40500